# DR2Seg: Decomposed Two-Stage Rollouts for Efficient Reasoning Segmentation in Multimodal Large Language Models

### v1

@@ Abstract
Reasoning segmentation is an emerging vision-language task that requires reasoning over intricate text queries to precisely segment objects. However, existing methods typically suffer from overthinking, generating verbose reasoning chains that interfere with object localization in multimodal large language models (MLLMs). To address this issue, we propose DR2Seg, a self-rewarding framework that improves both reasoning efficiency and segmentation accuracy without requiring extra thinking supervision. DR2Seg employs a two-stage rollout strategy that decomposes reasoning segmentation into multimodal reasoning and referring segmentation. In the first stage, the model generates a self-contained description that explicitly specifies the target object. In the second stage, this description replaces the original complex query to verify its self-containment. Based on this design, two self-rewards are introduced to strengthen goal-oriented reasoning and suppress redundant thinking. Extensive experiments across MLLMs of varying scales and segmentation models demonstrate that DR2Seg consistently improves reasoning efficiency and overall segmentation performance.
@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; Bai et al., 2025; OpenAI., 2024) encode rich open-world knowledge and exhibit strong capabilities in joint image-text understanding across downstream tasks. By leveraging these strengths, MLLMs can analyze visual scenes and interpret ambiguous human intents, enabling more complex tasks. Reasoning segmentation (Yu et al., 2016; Lai et al., 2024; Zhu et al., 2025) builds upon MLLMs by emphasizing the synergy between reasoning and perception, aiming to infer target objects from complex textual queries and achieve precise segmentation. Compared to the referring segmentation task (Ding et al., 2021; Yang et al., 2022; Liu et al., 2023c), reasoning segmentation involves more intricate and implicit queries, making it more challenging and more suitable for real-world agent scenarios, such as interactive robotics (Yin et al., 2023) and autonomous driving (Tian et al., 2024).
@@
Currently, there are two main paradigms in reasoning segmentation. Supervised fine-tuning (SFT)-based methods (Lai et al., 2024; Ren et al., 2024), following the pioneering work LISA (Lai et al., 2024), integrate pre-trained MLLMs with segmentation models through additional supervised fine-tuning to enable reasoning segmentation. While effective on in-domain tasks, SFT-based methods exhibit limited generalization to out-of-distribution (OOD) scenarios and lack explicit reasoning chains for explainability (Liu et al., 2025a; Wang et al., 2025). In contrast, reinforcement learning (RL)-based methods (Liu et al., 2025b; Wang et al., 2025), inspired by ZeroSeg (Liu et al., 2025a), optimize MLLMs with perception-oriented rewards using group-relative policy optimization (GRPO) (Shao et al., 2024). These methods can adaptively generate reasoning chains, achieving improved OOD generalization, which has attracted increasing attention. However, existing RL-based methods often suffer from overthinking, generating verbose reasoning chains that not only reduce computational efficiency but also interfere with accurate object localization. Recently, PixelThink (Wang et al., 2025) addresses this issue by estimating problem difficulty, but at the cost of introducing an additional large-scale expert MLLM for supervision (e.g., Qwen2.5-VL-72B (Bai et al., 2025)).
In this paper, we propose DR2Seg, a self-rewarding framework that achieves efficient reasoning and accurate localization without relying on extra supervision from other MLLMs. Specifically, we propose a two-stage rollout strategy that decomposes reasoning segmentation into two stages: multimodal reasoning and referring segmentation. Training involves two rollout passes of the same MLLM. (1) First pass: where the model generates an explicit inferring description to specify the target objects. (2) Second pass: the model is then re-prompted to reason based on the explicit description. If the derived answer is correct, the inferring description is regarded as faithful and receives a self-reward. This two-stage rollout improves the alignment between reasoning and description for accurate localization while reducing redundant reasoning. Moreover, we introduce a length-based self-reward that encourages more compact reasoning in the second stage under explicit description guidance, thereby further reducing thinking token counts. Our contributions are summarized as:
@@
  • DR2Seg designs a two-stage rollout strategy that decouples multimodal reasoning and perception in MLLM for reasoning segmentation, combined with a length-based self-reward to reduce redundant reasoning.
@@ 2.1 Reasoning Segmentation
Image segmentation has evolved from the traditional closed-set setting (Ronneberger et al., 2015; Minaee et al., 2021) to open-vocabulary setting of referring segmentation (Ding et al., 2021; Yang et al., 2022; Liu et al., 2023c), where objects are segmented using brief descriptions.
Recently, benefiting from the visual-language reasoning capabilities of MLLMs, reasoning segmentation (Shen et al., 2025) has further expanded prompts from fixed vocabularies to arbitrary linguistic forms. The pioneering work LISA (Shen et al., 2025) integrates MLLMs with the segmentation model SAM (Kirillov et al., 2023) by aligning textual reasoning with segmentation. Building on LISA, a series of studies explore supervised fine-tuning to strengthen the alignment between textual tokens and fine-grained segmentation (Yang et al., 2023; Ren et al., 2024). However, SFT-based methods suffer from limited generalization, leading to notable performance degradation in OOD scenarios. Seg-Zero (Liu et al., 2025a) addresses this issue by introducing an reinforcement learning based framework, which leverages GRPO (Shao et al., 2024) to adaptively optimize the model’s reasoning ability, achieving improved generalization. VisionReasoner (Liu et al., 2025b) further extends to enable multi-object segmentation by incorporating a bipartite matching algorithm. PixelThink (Wang et al., 2025) focuses on the efficiency and introduces an auxiliary large-scale MLLM to estimate query difficulty, thereby improving reasoning efficiency. In contrast, this paper proposes a self-reward framework that requires no additional MLLMs while achieving more efficient and accurate performance.
@@ 2.2 Self-Rewarding Reinforcement Learning
High-quality rewards are critical for reinforcement learning with verifiable rewards (RLVR), which typically relies on high-quality reward models or even human feedback, becoming a major bottleneck for scalability (Peng et al., 2025; Wen et al., 2025; Su et al., 2025). To address this dilemma, recent works have explored self-rewarding approaches, where reward signals are derived from the model itself (Yuan et al., 2024). In language-model-based settings, self-rewarding methods replace external reward models with signals such as model confidence (Li et al., 2025a; van Niekerk et al., 2025), uncertainty (Zhao et al., 2025), or self-verified solutions (Simonds et al., ).
Recently, several studies have extended this paradigm to MLLMs. For example, Calibrated Self-Rewarding (Zhou et al., 2024) iteratively generates responses and performs self-scoring, assigning rewards through progressively applied visual constraints. PLARE (Luu et al., 2025) queries a MLLM to obtain preference labels over pairs of visual trajectory segments, and directly trains the policy with a supervised contrastive preference learning, eliminating the need for an explicit reward model. Vision-SR1 (Li et al., 2025b) decomposes MLLM reasoning into visual perception and language reasoning by explicitly rewarding visual perception by the MLLM itself. This work further explores self-rewarding for reasoning perception in MLLMs, enabling concise reasoning chains that explicitly specify targets.
@@ 3.1 Problem Definition
Given an input image $\mathcal{I}$ and an implicit textual query $\mathcal{Q}$, reasoning segmentation (Lai et al., 2024) aims to produce a binary segmentation mask $\mathcal{M}$. This task is similar to referring segmentation (Ding et al., 2021) but is more challenging, as reasoning segmentation involves more complex queries expressed in arbitrary free-form natural language. Moreover, reasoning segmentation also emphasizes the generation of reasoning chains $\mathcal{R}$, which play a crucial role in understanding intent and reasoning to identify targets, thereby improving interpretability and generalization.
@@ 3.2 Overview
We follow the standard reasoning segmentation framework (Liu et al., 2025b) as shown in Fig. 2(a), which includes an reasoning MLLM and a segmentation model. The MLLM takes an image $\mathcal{I}$ and a textual query $\mathcal{Q}$ as input, and produces two outputs: $\mathcal{R},\mathcal{A}=\text{MLLM}(\mathcal{I},\mathcal{Q})$. Here, $\mathcal{R}$ denotes reasoning chains of the MLLM and $\mathcal{A}$ represents spatial answers, including a bounding box and a point, which serve as inputs of the segmentation model. We adopt the SAM series model (Kirillov et al., 2023; Ravi et al., ) as the segmentation model, which takes the image $\mathcal{I}$ and the spatial answers $\mathcal{A}$ as input and generate the binary masks $\mathcal{M}$ of targets.
@@ 3.3 DR2Seg: A Pure Self-Reward Framework
As discussed, redundant over-thinking in MLLMs can stem from attending to excessive objects, which confuses subsequent localization and degrade both efficiency and accuracy. However, supervising the reasoning process of MLLMs is inherently challenging because multiple reasoning paths can lead to correct answers, which largely explains the limited generalization of SFT-based methods. In contrast, RL-based approaches typically leave reasoning unconstrained, making them prone to over-thinking. PixelThink mitigates this issue by introducing an additional expert MLLM to constrain reasoning length, but this implicitly injects external knowledge from a larger model (i.e., Qwen2.5VL-72B), raising concerns about fairness and limiting the model’s self-evolution capability. To address these issues, we propose a self-reward framework consisting of two-stage rollout and self-reward design, as illustrated in Fig. 2(b) and Fig. 2(c).
Two-Stage Rollout. To encourage MLLMs to produce self-contained reasoning for segmentation, we enforce a think–description generation format. Given an image $\mathcal{I}$ and a textual query $\mathcal{Q}$, the MLLM outputs a structured response as: <think> $\mathcal{R}$ </think><description> $\mathcal{D}$ </description> <answer> $\mathcal{A}$ </answer>, where $\mathcal{R}$ is reasoning chains, $\mathcal{D}$ is inferring descriptions, and $\mathcal{A}$ is spatial answers for segmentation.
@@
Self-Reward Design. We can combine the two rollout passes into a longer reasoning path, formalized as: ($\mathcal{I}$, $\mathcal{Q}$) $\rightarrow$ $\mathcal{R}^{1}$ $\rightarrow$ $\mathcal{D}$ $\rightarrow$ ($\mathcal{I}$, $\mathcal{D}$) $\rightarrow$ $\mathcal{R}^{2}$ $\rightarrow$ $\mathcal{A}^{2}$. Based on this reasoning path, we infer whether the preceding input is self-contained according to the correctness of the answer. We then derive two guiding principles and design corresponding self-rewards:
@@
Accurate segmentation heavily relies on language descriptions, given the diverse number and granularity of objects in images. Therefore, the answer from the second rollout pass serves to verify the reasoning and informational completeness of the first-pass generation. If the model can still produce the correct answer given only $(I,\mathcal{D})$, we consider $\mathcal{D}$ to be correct and faithful, and accordingly assign a referring description self-reward $\textbf{R}_{\text{desc}}$:
@@
$\textbf{R}_{\text{desc}}(\mathcal{I},\mathcal{D})=\mathbb{I}[\mathcal{A}^{2}=\mathcal{A}^{*}],$ (2)
where $\mathcal{A}^{*}$ is the ground-truth answer.
@@
After multimodal reasoning in the first rollout pass, $\mathcal{D}$ serves as a concise phrase-level description of the target. Consequently, the language query in the second pass is simplified, leading to shorter reasoning chains $\mathcal{R}^{2}$ compared to those in the first pass $\mathcal{R}^{1}$. We first compute the token number of $\mathcal{R}^{1}$ and $\mathcal{R}^{2}$:
@@
Here, the first term $\mathbb{I}\!\left[\mathcal{N}^{2}<\mathcal{N}^{1}\right]$ follows Principle 2 by comparing the reasoning lengths of the two rollout passes, thereby promoting more concise and refined reasoning. However, since the first term is a purely comparative reward, using it alone lacks an absolute anchor and allows the model to exploit reward hacking: $\mathcal{N}^{1}$ and $\mathcal{N}^{2}$ increase synchronously, resulting in an overall growth of reasoning length. Therefore, we introduce the second term: $\gamma\,\max\!\big(0,\,\mathcal{N}^{1}-\mathcal{N}_{0}\big)$, where $\mathcal{N}_{0}$ is a predefined length anchor and $\gamma$ controls the strength of the length penalty. Finally, we apply a clip($\cdot$) operator to ensure that $\mathbf{R}_{\text{len}}$ remains within the range [0,1].
Instead of relying on an external reward model (e.g., Qwen2.5VL-72B in PixelThink), we leverage the model’s own capabilities for self-evaluation. By verifying the correctness of the second-pass answer and comparing the reasoning lengths across the two passes, our method enables self-rewarding over the reasoning process.
Total Reward. The total reward comprises three components, each conditioned on the image-query pair ($\mathcal{I}$, $\mathcal{Q}$).
Base reward $\mathbf{R}_{\text{base}}$: We adopt the base reward design from VisionReasoner, which consists of a format reward, a non-repeart reward, and a segmentation accuracy reward:
$\mathbf{R}_{\text{base}}$ $=\mathbf{R}_{\text{format}}+\mathbf{R}_{\text{non-repeat}}+\mathbf{R}_{\text{acc}},$ (6)
which together encourage structured outputs, discourage repetitive responses, and ensure accurate segmentation.
The total reward is computed as:
@@
where $\mathbf{R}_{\text{desc}}$ is the description self-reward, which evaluates the correctness of the answer generated in the second rollout pass. $\mathbf{R}_{\text{len}}$ is the length-based self-reward, which self-evaluates the difficult of question, ensuring that $\mathcal{D}$ is easier than $\mathcal{Q}$. $\tilde{\mathbf{R}}_{\text{len}}$ is the conditional length-based self-reward, defined as:
@@
When all $n$ rollouts yield zero accuracy reward, the length reward is disabled by setting $\mathbf{R}_{\text{len}}=1$, preventing premature length constraints before successful target localization. Inspire by  (Wang et al., 2025), we integrate the length-based reward through a multiplicative combination to ensure stability during training.
Training with GRPO. We adopt Group-Relative Policy Optimization (GRPO) to fine-tune the model via reinforcement learning, maximizing the total reward $\mathbf{R}_{\text{total}}$ in Eq. 7 for reasoning quality, segmentation accuracy, and token efficiency. By evaluating rewards at the mini-batch group level, GRPO reduces gradient variance, stabilizes training, and accelerates convergence.

### v2

@@ Abstract
Reasoning segmentation is an emerging vision-language task that requires reasoning over intricate text queries to precisely segment objects. However, existing methods typically suffer from overthinking, generating verbose reasoning chains that interfere with object localization in multimodal large language models (MLLMs). To address this issue, we propose DR2Seg, a self-rewarding framework that improves both reasoning efficiency and segmentation accuracy without requiring extra thinking supervision. DR2Seg employs a two-stage rollout strategy that decomposes reasoning segmentation into multimodal reasoning and referring segmentation. In the first stage, the model generates a self-contained description that explicitly specifies the target object. In the second stage, this description replaces the original complex query to verify its self-containment. Based on this design, two self-rewards are introduced to mitigate overthinking and the associated attention dispersion. Extensive experiments conducted on 3B and 7B variants of Qwen2.5-VL, as well as on both SAM2 and SAM3, demonstrate that DR2Seg consistently improves reasoning efficiency and overall segmentation accuracy.
@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; Bai et al., 2025; OpenAI., 2024) encode rich open-world knowledge and exhibit strong capabilities in joint image-text understanding across downstream tasks. By leveraging these strengths, MLLMs can analyze visual scenes and interpret ambiguous human intents, enabling more complex tasks. Recently, reasoning segmentation (Yu et al., 2016; Lai et al., 2024; Zhu et al., 2025) which builds upon MLLMs and emphasizes the synergy between reasoning and perception, has garnered widespread attention. Its core objective is to segment target objects from complex textual queries. Compared to the referring segmentation task (Ding et al., 2021; Yang et al., 2022; Liu et al., 2023c), reasoning segmentation involves more intricate and implicit queries, making it more challenging and more suitable for real-world agent scenarios, such as interactive robotics (Yin et al., 2023) and autonomous driving (Tian et al., 2024).
@@
Currently, there are two main paradigms in reasoning segmentation. Supervised fine-tuning (SFT)-based methods (Lai et al., 2024; Ren et al., 2024), which integrate pre-trained MLLMs with segmentation models through SFT to enable reasoning segmentation. However, SFT-based methods exhibit limited generalization to out-of-distribution (OOD) scenarios and lack explicit reasoning chains for explainability (Liu et al., 2025a; Wang et al., 2025). In contrast, reinforcement learning (RL)-based methods (Liu et al., 2025b; Wang et al., 2025), inspired by ZeroSeg (Liu et al., 2025a), optimize MLLMs with perception-oriented rewards using GRPO (Shao et al., 2024). By adaptively generating reasoning chains, these methods can achieve improved OOD generalization. However, existing RL-based methods often suffer from overthinking, generating verbose reasoning chains that not only reduce computational efficiency but also interfere with accurate object localization. To address this issue, PixelThink (Wang et al., 2025) leverages an extra large-scale expert MLLM (i.e., Qwen2.5-VL-72B (Bai et al., 2025)) to estimate problem difficulty, thereby implicitly introducing thinking supervision. However, such a strategy depends strongly on the expert MLLM as an auxiliary module, without probing into the intrinsic self-organizing reasoning capacities of the base model itself.
These issues motivate us to explore self-rewarding (Yuan et al., 2024; Zhou et al., 2024), a recent advancement in reasoning MLLMs that remains underexplored in reasoning perception. Unlike visual question answering in reasoning MLLMs, reasoning perception is more prone to attention confusion due to inherent modality differences between coordinates and text. To address this challenge, we propose DR2Seg, which adopts a two-stage rollout strategy to decompose reasoning segmentation into multimodal reasoning and referring segmentation. Training involves two rollout passes of the same MLLM: 1) In the first pass, the model generates an explicit inferring description to specify target objects. 2) In the second pass, it is re-prompted to respond based on this description. If the second-pass prediction is correct, the description is regarded as faithful and receives a self-reward. Furthermore, we introduce a length-based self-reward that encourages concise reasoning under explicit description guidance, thereby reducing unnecessary thinking tokens. As a result, DR2Seg achieves efficient reasoning and accurate localization without thinking supervision, as shown in Fig. 1. Our contributions are summarized as:
@@
  • DR2Seg designs a two-stage rollout strategy that decouples multimodal reasoning and perception in MLLM for accurate segmentation, combined with a length-based self-reward to reduce redundant reasoning.
@@ 2.1 Reasoning Segmentation
Image segmentation has evolved from the traditional closed-set setting (Ronneberger et al., 2015; Minaee et al., 2021) to open-vocabulary setting of referring segmentation (Ding et al., 2021; Yang et al., 2022; Liu et al., 2023c), where objects are segmented using brief descriptions. Recently, benefiting from the visual-language reasoning capabilities of MLLMs, reasoning segmentation (Shen et al., 2025; Zhu et al., 2026) has further expanded prompts from fixed vocabularies to arbitrary linguistic forms.
The pioneering work LISA (Shen et al., 2025) integrates MLLMs with the segmentation model SAM (Kirillov et al., 2023) by aligning textual reasoning with segmentation. Building on LISA, a series of studies explore supervised fine-tuning to strengthen the alignment between textual tokens and fine-grained segmentation (Yang et al., 2023; Ren et al., 2024). However, SFT-based methods suffer from limited generalization, leading to notable performance degradation in OOD scenarios. Seg-Zero (Liu et al., 2025a) addresses this issue by introducing an reinforcement learning based framework, which leverages GRPO (Shao et al., 2024) to adaptively optimize the model’s reasoning ability, achieving improved generalization. VisionReasoner (Liu et al., 2025b) further extends to enable multi-object segmentation by incorporating a bipartite matching algorithm. PixelThink (Wang et al., 2025) focuses on the efficiency and introduces an auxiliary large-scale MLLM to estimate query difficulty, thereby improving reasoning efficiency. In contrast, this paper proposes a self-reward framework that requires no additional MLLMs while achieving more efficient and accurate performance.
@@ 2.2 Self-Rewarding Reinforcement Learning
High-quality rewards are critical for reinforcement learning with verifiable rewards (RLVR), which typically relies on high-quality reward models or even human feedback, becoming a major bottleneck for scalability (Peng et al., 2025; Wen et al., 2025; Su et al., 2025). To address this dilemma, recent works have explored self-rewarding approaches, where reward signals are derived from the model itself (Yuan et al., 2024). In language-model-based settings, self-rewarding methods replace external reward models with signals such as model confidence (Li et al., 2025a; van Niekerk et al., 2025), uncertainty (Zhao et al., 2025), or self-verified solutions (Simonds et al., 2025).
Recently, several studies have extended this paradigm to MLLMs. For example, Calibrated Self-Rewarding (Zhou et al., 2024) iteratively generates responses and performs self-scoring, assigning rewards through progressively applied visual constraints. PLARE (Luu et al., 2025) queries a MLLM to obtain preference labels over pairs of visual trajectory segments, and directly trains the policy with a supervised contrastive preference learning, eliminating the need for an explicit reward model. Vision-SR1 (Li et al., 2025b) decomposes MLLM reasoning into visual perception and language reasoning by explicitly rewarding visual perception. This work further explores self-rewarding for reasoning perception, focusing on instance-level object understanding through decomposition into reasoning and referring segmentation via an explicit description bottleneck.
@@ 3.1 Problem Definition
Given an input image $\mathcal{I}$ and an implicit textual query $\mathcal{Q}$, reasoning segmentation (Lai et al., 2024) aims to produce a binary segmentation mask $\mathcal{M}$. This task is similar to referring segmentation (Ding et al., 2021) but is more challenging, as reasoning segmentation involves more complex queries expressed in arbitrary free-form natural language. Moreover, reasoning segmentation also emphasizes the generation of reasoning chains $\mathcal{R}$, which play a crucial role in understanding intent and reasoning to identify targets, thereby improving explainability and generalization.
@@ 3.2 Overview
We follow the standard reasoning segmentation framework (Liu et al., 2025b) as shown in Fig. 2(a), which includes a reasoning MLLM and a segmentation model. The MLLM takes an image $\mathcal{I}$ and a textual query $\mathcal{Q}$ as input, and produces two outputs: $\mathcal{R},\mathcal{A}=\text{MLLM}(\mathcal{I},\mathcal{Q})$. Here, $\mathcal{R}$ denotes reasoning chains of the MLLM and $\mathcal{A}$ represents spatial answers, including a bounding box, a point, and an optional description, which serve as inputs of the segmentation model. We adopt the SAM series (Kirillov et al., 2023; Ravi et al., ; Carion et al., 2025) as the segmentation models, which take the image $\mathcal{I}$ and the answers $\mathcal{A}$ as input and generate the binary masks $\mathcal{M}$ for target objects.
@@ 3.3 DR2Seg: A Pure Self-Reward Framework
As discussed, redundant over-thinking in MLLMs confuses subsequent localization and degrades both efficiency and accuracy. However, supervising the reasoning process is inherently challenging because multiple reasoning paths can lead to correct answers, which largely explains the limited generalization of SFT-based methods. PixelThink (Wang et al., 2025) addresses this issue by introducing an expert MLLM to constrain reasoning length. However, this implicitly injects external knowledge from a larger model (i.e., Qwen2.5VL-72B), raising fairness concerns and hindering the model’s ability to self-evolve. Instead, we propose a self-reward framework consisting of two-stage rollout and self-reward design, as shown in Fig. 2(b) and Fig. 2(c).
Two-Stage Rollout. To encourage MLLMs to produce self-contained reasoning for segmentation, we enforce a “think then description” generation format. Given an image $\mathcal{I}$ and a textual query $\mathcal{Q}$, the MLLM outputs a structured response as: <think> $\mathcal{R}$ </think><description> $\mathcal{D}$ </description> <answer> $\mathcal{A}$ </answer>, where $\mathcal{D}$ denotes inferring descriptions.
@@
Self-Reward Design. We can combine the two rollout passes into a longer reasoning path based on the token generation order of the MLLM, formalized as: ($\mathcal{I}$, $\mathcal{Q}$) $\rightarrow$ $\mathcal{R}^{1}$ $\rightarrow$ $\mathcal{D}$ $\rightarrow$ ($\mathcal{I}$, $\mathcal{D}$) $\rightarrow$ $\mathcal{R}^{2}$ $\rightarrow$ $\mathcal{A}^{2}$. Based on this reasoning path, we infer whether the preceding input is self-contained according to the correctness of the answer. We then derive two guiding principles and design corresponding self-rewards:
@@
Accurate segmentation heavily relies on language descriptions, given the diverse number and granularity of objects in the image. Therefore, the answer from the second rollout pass serves to verify the reasoning and informational completeness of the first-pass generation. If the model can still produce the correct answer given only $(I,\mathcal{D})$, we consider $\mathcal{D}$ to be correct and faithful, and accordingly assign a description self-reward $\textbf{R}_{\text{desc}}$:
@@
$\textbf{R}_{\text{desc}}(\mathcal{A}^{2},\mathcal{A}^{*})=\textbf{R}_{\text{acc}}(\mathcal{A}^{2},\mathcal{A}^{*}),$ (2)
where $\mathcal{A}^{*}$ is the ground-truth answer and $\textbf{R}_{\text{acc}}$ is the answer accuracy reward as in (Liu et al., 2025b).
@@
After multimodal reasoning in the first pass, $\mathcal{D}$ provides a concise phrase-level description of the target, which should lead to shorter reasoning chains $\mathcal{R}^{2}$ compared to those in the first pass $\mathcal{R}^{1}$. We first compute the token number of $\mathcal{R}^{1}$ and $\mathcal{R}^{2}$:
@@
Here, the first term $\mathbb{I}\!\left[\mathcal{N}^{2}<\mathcal{N}^{1}\right]$ follows Principle 2 by comparing the reasoning lengths of the two passes, thereby promoting more concise and refined reasoning. However, since the first term is a purely comparative reward, using it alone lacks an absolute anchor and allows the model to exploit reward hacking: $\mathcal{N}^{1}$ and $\mathcal{N}^{2}$ increase synchronously. Therefore, we introduce the second term: $\gamma\,\max\!\big(0,\,\mathcal{N}^{1}-\mathcal{N}_{0}\big)$, where $\mathcal{N}_{0}$ is a predefined length anchor and $\gamma$ controls the strength of the length penalty. Finally, we apply a clip($\cdot$) operator to ensure that $\mathbf{R}_{\text{len}}$ remains within the range [0,1].
Instead of relying on an external reward model (e.g., Qwen2.5VL-72B in PixelThink (Wang et al., 2025)), we leverage the model’s own capabilities for self-evaluation. By verifying the correctness of the second-pass answer and comparing the reasoning lengths across the two passes, our method enables self-rewarding over the reasoning process.
Total Reward. First, we adopt the base reward $\mathbf{R}_{\text{base}}$ from the baseline VisionReasoner (Liu et al., 2025b), which consists of a format reward, a non-repeat reward, and a answer accuracy reward:
$\mathbf{R}_{\text{base}}$ $=\mathbf{R}_{\text{format}}+\mathbf{R}_{\text{non-repeat}}+\mathbf{R}_{\text{acc}}.$ (6)
These rewards encourage structured output, discourage repetitive responses, and ensure accurate segmentation.
Then, we incorporate the proposed description self-reward $\mathbf{R}_{\text{desc}}$ and length-based self-reward $\mathbf{R}_{\text{len}}$ into the overall reward. The total reward is computed as:
@@
$\mathbf{R}_{\text{desc}}$ evaluates the correctness of the answer generated in the second rollout pass. $\tilde{\mathbf{R}}_{\text{len}}$ denotes the conditional length-based self-reward, defined as:
@@
When all $n$ rollouts yield zero accuracy reward, $\tilde{\mathbf{R}}_{\text{len}}$ is disabled by setting to 1. This avoids imposing premature length constraints before successful target localization and encourages more active exploration in the early stage. In this way, the model can dynamically balance reasoning length against answer accuracy based on problem difficulty. To ensure training stability, we integrate $\tilde{\mathbf{R}}_{\text{len}}$ into the reward function using a multiplicative formulation.
Training with GRPO. We adopt Group-Relative Policy Optimization (GRPO) (Shao et al., 2024) to fine-tune the model, maximizing the total reward $\mathbf{R}_{\text{total}}$ in Eq. 7 for reasoning quality, segmentation accuracy, and token efficiency. By evaluating rewards at the mini-batch group level, GRPO avoids the use of a reward model and stabilizes training. More details are provided in Sec. A of the Appendix.